Determine the pH during the titration of 73.1 mL of 0.462 M benzoic acid (Ka = 6.3×10-5) by 0.462 M NaOH at the following points. (a) Before the addition of any NaOH (b) After the addition of 17.0 mL of NaOH (c) At the half-equivalence point (the titration midpoint) (d) At the equivalence point (e) After the addition of 110 mL of NaOH
Answer #1

Phwol c o с (-) T Ph wo- tit o и mult where is the mitral concentration & be the degree of democrationka= - 1 Pho) HD boh woh ca.ca car an del :1- - Снт - се = c skape =skat - Light] = L Kagkat & logeentaoraltosess of [ht] =kaxc lorarea 16.3x10° X 0.462 tra protest - pH = -log THT) horfurst orto = wg ($6.3X105X0-462) 2.27 ©[ PhWOH) = (3-1-17) X0.462 (73.1+17) -0. 288 m (Phcoona] pH = pra tvog (phcoot) oost =-wg (6.3x105) +log - 0 288 -3.68 c) At1o) at eqr. these contarm only Phevona. NaCo HS Co, or Colts coo wa is a salt of weak aust (Phcoolt) and strong base (araoh P18 km Again Ghcoot COH phcooT (Phwoh Cot cut phwo-411 Now akn = - Ph wot phwo- + ht Ka - tphcooth t Phloot). H+ Cow - ON PhcoAt equ. pont VNaox Saн — Урилоон X SpoDo2 - NNOoH = 0.4624 73.1 0.462 = 13-1 -- [phiona) = _ 73.1X0.462 (73.17 173.1) = 0.273who, 100 ml NOW H added. the son contain 110731 NuoH . , 36.9 ml of exam m (OH) = 360nly 0.462 (110+73.1 mm is 0.093m thenpH